lamar's summer job is mowing lawns for landscapers. his pay is $7.50/h. Lamar also makes $11.25/h for any time over 40h that he works in one week. he worked 40 h last week plus n overtime hours and made $339.38. how much over time hours did he work?
Answer by checkley79(3341)

7.50*40+11.25X=339.38
300+11.25X=339.38
11.25X=339.38-300
11.25X=39.38
X=39.38/11.25
X=3.5 HOURS OF OVERTIME.